2. For a conjugated molecule with 18 π electrons.
a. Calculate the length of the molecules
b. Write the value of n for the highest occupied level
c. Calculate the longest wavelength transition for this molecule using the 1 dimensional

I have an exam coming up soon and I am having trouble understanding these practice questions.

particle in a box model.
Answer #1

1877 electrons conjugaled molecule HC CH=CH-CH=CH-CH=CH-CH=CH-CH=CH-CH=CH. HC HC = CH H = HC-HC length of the molecule calculd = 27,55 A°= 27,55×100m h = 6.625x1034 TS electronmays me = 9:llx LE = (6:625 X1734 IS) 8xquxő/kg *(27,95x160m) ng *(7.95 x